Calculate Kb of Mg(OH)2 . It has a pH of 9.65 and the concentration is 0.10M

Mg(OH)2+ H2O----> Mg2++ 2OH-
Answer #1

Mogte + 2014 ty O + Mg (OH) - 0.40M 2(0:1) = 0.2M concentration of oth = o.am pH = 9.65 ; POH= 14-pH poH = 14–9.65 poll = 4.3log K = -8.001 Man = antilog (-8.001) Kg = 108001108,-0.001 Ab = 0.998108 Wag = 1x168 Scanned with CamScanner